OpenStudy (anonymous):

For a cone, if the height is 5mm and the radius is 3mm then the exact volume is....?

OpenStudy (phi):

Volume of a cone is V= B*H/3 See http://en.wikipedia.org/wiki/Cone_(geometry)#Volume where B is the area of the base, and H is the height

OpenStudy (phi):

So you need to be able to find the area of the base (a circle with radius 3 mm) the multiply by the height. then divide by 3

OpenStudy (pfenn1):

@phi is correct, I read the shape as a cylinder, not as a cone.

OpenStudy (phi):

Area of a circle is \( \pi r^2 \) where r= radius. r= 3 mm for this problem